# Observational predictions of some inflationary models

### v3

@@ 4 The generalized Starobinsky inflationary model
The generalized Starobinsky inflationary model is given by [16, 33], and it is a slight modification of the Starobinsky inflationary model
@@
$V_{0}=6\left(\dfrac{2p-1}{4p}\right)M^{2}\left(\dfrac{1}{2p}\right)^{\dfrac{1}{2p-1}}.$ (20)
@@
From Eqs. (4)–(7) using the potential given by Eq. (19), we calculate using Wolfram Mathematica® the slow–roll parameters, and following the notation of Renzi [16] we obtain that
@@
$\omega_{\nu}$ $=$ $\dfrac{64}{27\left(2p-1\right)^{6}\left(F_{1}-1\right)^{6}}\Bigg{[}\left(512p^{6}-1536p^{5}+1920p^{4}-1280p^{3}+480p^{2}-96p+8\right)\Bigg{.}$
@@
Fig. 3 displays the CMB angular spectrum reproduced in CAMB for various values of the parameter $p$ in comparison to Planck $2018$ data. It is evident that for $p>1$, this parameter shifts the spectrum upwards, whereas for $p<1$ it shifts it downwards in comparison to the Starobinsky inflationary model ($p=1$). As the reproduced CMB by the Starobinsky inflationary model was under the observational data, we expect the optimal value of $p$ to be a little bit greater than $1$ to make it approach the observational curve.
Performing the CAMB reproduction of the CMB for various values of $p$, we found that the value that gives us the best tensor–to–scalar ratio $r$ and the scalar spectral index $n_{\mathrm{S}}$, in comparison with the ones reported by Planck $2018$ results, is $1.0004\pm 0.0001$, which is greater than $1$ as predicted before. Now let us plot it.
@@
When considering all scales, this inflationary model has an L2222 norm error of $43$ $\mu K^{2}$. At large, intermediate, and small angular scales, the L2222 norm errors are $34$ $\mu K^{2}$, $65$ $\mu K^{2}$, and $22$ $\mu K^{2}$, respectively. Therefore, similar to the Starobinsky inflationary model, the generalized Starobinsky inflationary model reproduces the observational CMB angular spectrum more accurately at smaller angular scales, but the error balance is less favorable at intermediate scales. However, it significantly outperforms the Starobinsky inflationary model both in terms of relative error size and error balance
@@ 4.1 The Acoustic Peaks
In Table 3, we present the acoustic peaks identified in the intermediate range of the spectrum, along with the corresponding troughs. Our observations reveal close proximity in the positions along the multipole axis between the observed and calculated values. Considering the uncertainties associated with the observed data, all computed peaks, with the exception of the third one, align well with the Planck satellite. The percentage errors associated with the calculated multipole values for peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $0.18$, $1,27$, $0.20$, $0.22$, and $0.52$, respectively. This errors are exactly the same of those reported by the Starobinsky inflationary model.
@@
4.2 Cosmic Parameters
In Table 4, we compare certain cosmic parameters computed theoretically using the generalized Starobinsky inflationary model with their corresponding observational values derived from the Planck satellite. Once again, all parameters demonstrate alignment between their theoretical predictions and observed values. In fact, all values remain the same as those derived from the Starobisky model ($p=1$), except for the scalar spectral index, which has a smaller uncertainty this time.
Cosmic Paramenter Symbol G. Starobinsky Planck
@@ 4.2 Cosmic Parameters
Table 4: Cosmic parameters obtained from the generalized Starobinsky inflationary model with $p=1.0004$ compared with the cosmic parameters reported by Planck $2018$ results.
The percentage errors associated with the age of the universe, matter density, baryon density, dark energy density, and scalar spectral index are $0.007$, $0.16$, $0.04$, $0.09$, and $0.2$, respectively. This model exhibits the same capacity for accurate and precise predictions of the observed cosmic parameters as the Starobinsky inflationary model. However, due to the smaller uncertainty in the scalar spectral index, we can say that this generalized model is slightly better in that aspect, although the difference is not significant.
@@ 5 Chaotic inflationary model with a step
where the step occurs at $\phi=\phi_{\textnormal{step}}$, $\phi$ is the inflaton, $m$ is the inflaton mass, and the parameters $c$ and $d$ are related to the amplitude and width of the feature, respectively [18, 19, 20, 21, 22].
Using Eqs. (4)–(7) we calculated also using Wolfram Mathematica® the slow-roll parameters for the potential (25) of the chaotic inflationary model with a step, obtaining
@@
$\xi_{\nu}$ $=$ $\dfrac{2}{\phi^{3}}F_{2}^{2}\left[c\phi F_{2}^{2}+2\left(d+cdF_{3}\right)\right]\Bigg{[}\dfrac{\left(3d^{2}+2\phi^{2}-3\phi F_{2}^{2}\right)\left(\phi+dF_{4}\right)}{d^{4}\left(1+cF_{3}\right)^{2}}\Bigg{]},$ (28)
@@
Fig. 6 shows the CMB angular power spectrum for the chaotic inflationary model with a step, different values of the parameters $c$ and $d$ were used to reproduce different spectra using CAMB. In Fig. 7(a) we report the best fit compared to Planck $2018$ data, for this fit we use $d=0.04$ and $c=0.12$. Our best fit was chosen by computing the relative error between our model and Planck $2018$ data and selecting the parameters that produce the lower relative error. We can see from Fig. 8(a) that the relative error found for these parameters is lower than 9% concerning the spectrum from Planck $2018$ data.
Figure 6: Angular power spectrum for the chaotic inflationary model with a step reproduced by CAMB for different values of $c$ and $d$. Red line: Planck $2018$ data, brown line: result for $c=0$, turquoise line: result for $c=0.12$ and $d=0,04$, gray line: result for $c=0.012$ and $d=0,04$, and pink line: result for $c=0.0012$ and $d=0,04$.
@@
This model reproduces a CMB that aligns better with the Planck $2018$ observations, surpassing the Starobinsky and Generalized Starobinsky inflationary models. High–variability in the results persists at small angular scales as in the previous models. To assess the overall balance of these errors, let’s examine the L2222 norm.
When considering all scales, this inflationary model has a L2222 norm error of 26 $\mu K^{2}$. At large, intermediate, and small angular scales, the L2 norm errors are 20 $\mu K^{2}$, 36 $\mu K^{2}$, and 18 $\mu K^{2}$, respectively. Therefore, the chaotic model exhibits similar characteristics as the models seen until now. It reproduced the CMB angular spectrum more accurately at smaller angular scales, but the error balance is less favorable at intermediate scales. But, in general, it is the best model for reproducing the CMB angular spectrum, both in terms of relative error size and error balance.
@@ 5.1 The Acoustic Peaks
In Table 5, we present the acoustic peaks identified in the intermediate scale of the spectrum, along with the corresponding troughs in this range.
@@
Once again, our observations reveal proximity between the positions on the multipole axis of both observed and computed peaks. Taking into account the errors associated with the observed values, all computed peaks, except the third one, align well with the data from the Planck satellite. Specifically, the percentage errors for the calculated multipole values corresponding to peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $0.18$, $1.27$, $0.20$, $0.22$, and $0.52$, respectively. We can see again the same percentage error of the previously presented models.
@@
5.2 Cosmic Parameters
In Table 6, we once again analyze the same cosmic parameters derived theoretically from the chaotic model and compare them with their corresponding observational values obtained from the Planck satellite. Considering uncertainties, all the parameters exhibit alignment between their theoretical and observational values.
Cosmic Paramenter Symbol Chaotic Planck
@@ 5.2 Cosmic Parameters
Table 6: Cosmic parameters obtained from the chaotic inflationary model with a step for $p=1.0004$ compared with the cosmic parameters reported by Planck $2018$ results.
@@
The percentage errors for the age of the universe, matter density, baryon density, dark energy density, and scalar spectral index are $0.007$, $0.16$, $0.04$, $0.09$, and $0.2$, respectively. The effectiveness, accuracy, and precision demonstrated by the chaotic model are again similar to the previous ones. It can also be stated that it is slightly better, mainly due to the better precision it achieves when predicting the scalar spectral index.

### v4

@@ 4 The generalized Starobinsky inflationary model
The generalized Starobinsky inflationary model is given by [16, 34], and it is a slight modification of the Starobinsky inflationary model,
@@
$V_{0}=6\left(\dfrac{2p-1}{4p}\right)M^{2}\left(\dfrac{1}{2p}\right)^{\frac{1}{2p-1}}.$ (24)
@@
From Eqs. (4)–(7) using the potential given by Eq. (23), we calculate the slow–roll parameters using a symbolic manipulation program, and following the notation of Renzi [16] we obtain that:
@@
$\omega_{\nu}$ $=$ $\dfrac{64}{27\left(2p-1\right)^{6}\left(F_{1}-1\right)^{6}}\Bigg{[}\left(512p^{6}-1536p^{5}+1920p^{4}-1280p^{3}+480p^{2}-96p+8\right)\Bigg{.}$ (28)
@@
For this model, we need to calculate the values of $\phi_{\text{ini}}$ numerically using Eq. (8), and $\phi_{\text{end}}$ from Eq. (4) when $\epsilon_{\nu}=1$, which becomes:
$\phi_{\text{end}}=\sqrt{\dfrac{3}{2}}\ln\left[\dfrac{1-4p^{2}-2\sqrt{3\left(p^{2}-4p^{3}+4p^{4}\right)}}{1+4p^{2}-8p^{2}}\right],$ (29)
where Eq. (29) reduces to Eq. (21) for the Starobinsky inflationary model when $p=1$. In Fig. 5, we show the evolution of the scalar field $\phi(t)$ as a function of time $t$ of the generalized Starobinsky inflationary model considering $p=1.0004$. The figure illustrates the oscillation of the scalar field $\phi(t)$ after the end of inflation.
Figure 5: Evolution of the scalar field $\phi(t)$ as a function of time $t$ of the generalized Starobinsky inflationary model for $p=1.0004$.
Using Eq. (14) we obtain that the value for the scalar–to–tensor ratio $r$ for the generalized Starobinsky inflationary model for each considered case which is shown in Table 3. From table 3 we can observed that for $p=1$ we recover the result of the Starobinsky inflationary model.
$p$ $r$
0.995 0.00449759
0.996 0.00404992
0.997 0.00381766
0.998 0.00351217
1.000 0.0029639
1.0004 0.00286365
1.001 0.00271881
1.002 0.00249167
1.003 0.00228136
1.004 0.00208686
1.005 0.00190719
Table 3: Value of the scalar–to–tensor ratio $r$ for each value of $p$ considered in the generalized Starobinsky inflationary model.
Fig. 6(a) and 6(b) show that the slow–roll parameters (25) and (4) for the generalized Starobinsky inflationary model are satisfied.
Figure 6: Behaviour of the slow–roll parameters $\epsilon_{\nu}$ and $\eta_{\nu}$ with respect to the number of e–folds $N_{e}$, described by the generalized Starobinsky inflationary model for $p=1.0004$.
Fig. 7 shows the CMB angular spectrum reproduced in CAMB for multiple values of the parameter $p$ in comparison to Planck $2018$ data. It is evident that for $p>1$, this parameter shifts the spectrum upwards, whereas for $p<1$ it shifts it downwards in comparison to the Starobinsky inflationary model ($p=1$). As the CMB reproduced by the Starobinsky inflationary model was under the observational data, we expect the optimal value of $p$ to be a little bit greater than $1$ to approach the observational curve.
Performing the CAMB reproduction of the CMB for various values of $p$, we found that the value that gives us the best tensor–to–scalar ratio $r$ and the scalar spectral index $n_{\mathrm{S}}$, compared to the ones reported by Planck $2018$ results, is $1.0004\pm 0.0001$, which is greater than $1$ as predicted before. Now, let us plot it.
@@
When considering all scales, this inflationary model has an L2222 norm error of $43$ $\mu K^{2}$. At large, intermediate, and small angular scales, the L2222 norm errors are $34$ $\mu K^{2}$, $65$ $\mu K^{2}$, and $22$ $\mu K^{2}$, respectively. Therefore, similar to the Starobinsky inflationary model, the generalized Starobinsky inflationary model reproduces the observational CMB angular spectrum more accurately at smaller angular scales, but the error balance is less favorable at intermediate scales. However, it significantly outperforms the Starobinsky inflationary model in terms of both the relative error size and the error balance.
@@ 4.1 The Acoustic Peaks
In Table 4, we present the acoustic peaks identified in the intermediate range of the spectrum, together with the corresponding troughs. Our observations reveal close proximity in the positions along the multipole axis between the observed and calculated values. Considering the uncertainties associated with the observed data, all computed peaks, with the exception of the third one, align well with the Planck satellite. The percentage errors associated with the calculated multipole values for peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $0.18$, $1,27$, $0.20$, $0.22$, and $0.52$, respectively. This errors are exactly the same of those reported by the Starobinsky inflationary model.
@@
4.2 Cosmological parameters
In Table 5, we compare certain cosmological parameters computed theoretically using the generalized Starobinsky inflationary model with their corresponding observational values derived from the Planck satellite. Once again, all parameters demonstrate alignment between their theoretical predictions and observed values. In fact, all values remain the same as those derived from the Starobisky model ($p=1$), except for the scalar spectral index, which has a smaller uncertainty this time.
Cosmological Parameter Symbol G. Starobinsky Planck
@@ 4.2 Cosmological parameters
Table 5: Cosmological parameters obtained from the generalized Starobinsky inflationary model with $p=1.0004$ compared with the cosmological parameters reported by Planck $2018$ results.
The percentage errors associated with the age of the universe, matter density, baryon density, dark energy density, and scalar spectral index are $0.007$, $0.16$, $0.04$, $0.09$, and $0.2$, respectively. This model exhibits the same capacity for accurate and precise predictions of the observed cosmological parameters as the Starobinsky inflationary model. However, due to the smaller uncertainty in the scalar spectral index, we can say that this generalized model is slightly better in that aspect, although the difference is not significant.
@@ 5 Chaotic inflationary model with a step
where the step occurs at $\phi=\phi_{\textnormal{step}}$, $\phi$ is the inflaton, $m$ is the inflaton mass, and the parameters $c$ and $d$ are related to the amplitude and width of the feature, respectively [18, 19, 20, 36, 21, 22].
Using Eqs. (4)–(7) we calculated also, using a Software of symbolic manipulation, the slow–roll parameters for the potential (30) of the chaotic inflationary model with a step, obtaining
@@
$\xi_{\nu}$ $=$ $\dfrac{2}{\phi^{3}}F_{2}^{2}\left[c\phi F_{2}^{2}+2\left(d+cdF_{3}\right)\right]\Bigg{[}\dfrac{\left(3d^{2}+2\phi^{2}-3\phi F_{2}^{2}\right)\left(\phi+dF_{4}\right)}{d^{4}\left(1+cF_{3}\right)^{2}}\Bigg{]},$
@@
For this model, we need to calculate numerically the values of $\phi_{\text{ini}}$ using Eq. (8) and $\phi_{\text{end}}$ using Eq. (31) when $\epsilon_{\nu}=1$. In Fig. 10, we show the evolution of the scalar field $\phi(t)$ as a function of time $t$ of the chaotic inflationary model with a step for $c=0.12$, $d=0.04$, and $\phi_{\text{step}}=14.09$. The figure illustrates a step in the scalar field because the presence of the feature, and its oscillation after the end of inflation.
Figure 10: Evolution of the scalar field $\phi(t)$ as a function of time $t$ of the chaotic inflationary model with a step for $c=0.12$, $d=0.04$, and $\phi_{\text{step}}=14.09$.
Using Eq.(14) we obtain that the value for the scalar–to–tensor ratio $r$ for the chaotic inflationary model with a step for each considered case is,
$c$ $d$ $\phi_{\textnormal{step}}$ $r$
0 0.04 14.09 0.132232
0.0012 0.04 14.09 0.132232
0.012 0.04 14.09 0.132232
0.12 0.04 14.09 0.132232
Table 6: Value of the scalar–to–tensor ratio $r$ for each value of $c$, $d$, and $\phi_{\textnormal{step}}$ considered in the chaotic inflationary model with a step.
Fig. 11(a) and 11(b) show the slow–roll parameters (31) and (32) for the chaotic inflationary model with a step, in which we can observed that slow–roll parameters are violet at $N_{e}\sim 10$ just where the step occurs, it is because the presence of the feature [19, 20].
Figure 11: Behaviour of the slow–roll parameters $\epsilon_{\nu}$ and $\eta_{\nu}$ with respect to the number of e–folds $N_{e}$, described by the chaotic inflationary model with a step inflationary model for $c=0.12$, $d=0.04$, and $\phi_{\text{step}}=14.09$.
Fig. 12 shows the CMB angular power spectrum for the chaotic inflationary model with a step, different values of the parameters $c$ and $d$ were used to reproduce different spectra using CAMB. In Fig. 13(a) we report the best fit compared to Planck $2018$ data, for this fit we use $d=0.04$, $c=0.12$, and $\phi_{\text{step}}=14.09$. Our best fit was chosen by computing the relative error between our model and the Planck $2018$ data and selecting the parameters that produce the lower relative error. We can see from Fig. 14(a) that the relative error found for these parameters is lower than 9% regarding the spectrum from Planck $2018$ data.
Figure 12: Angular power spectrum for the chaotic inflationary model with a step reproduced by CAMB for different values of $c$ and $d$. Red line: Planck $2018$ data, brown line: result for $c=0$, turquoise line: result for $c=0.12$ and $d=0.04$, gray line: result for $c=0.012$ and $d=0.04$, and pink line: result for $c=0.0012$ and $d=0.04$.
@@
This model reproduces a CMB that aligns better with the Planck $2018$ observations, surpassing the Starobinsky and generalized Starobinsky inflationary models. High–variability in the results persists at small angular scales as in the previous models. To assess the overall balance of these errors, let us examine the L2222 norm.
When considering all scales, this inflationary model has a L2222 norm error of 26 $\mu K^{2}$. On large, intermediate and small angular scales, the L2 norm errors are 20 $\mu K^{2}$, 36 $\mu K^{2}$, and 18 $\mu K^{2}$, respectively. Therefore, the chaotic model exhibits similar characteristics as the models seen until now. It reproduced the CMB angular spectrum more accurately at smaller angular scales, but the error balance is less favorable at intermediate scales. But, in general, it is the best model for reproducing the CMB angular spectrum, both in terms of relative error size and error balance.
@@ 5.1 The Acoustic Peaks
In Table 7, we present the acoustic peaks identified in the intermediate scale of the spectrum, together with the corresponding troughs in this range.
@@
Once again, our observations reveal the proximity between the positions on the multipole axis of both observed and computed peaks. Taking into account the errors associated with the observed values, all computed peaks, except the third one, align well with the data from the Planck satellite. Specifically, the percentage errors for the calculated multipole values corresponding to peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $0.18$, $1.27$, $0.20$, $0.22$, and $0.52$, respectively. We can see again the same percentage error of the previously presented models.
@@
5.2 Cosmological parameters
In Table 8, we once again analyze the same cosmological parameters derived theoretically from the chaotic model and compare them with their corresponding observational values obtained from the Planck satellite. Considering uncertainties, all the parameters exhibit alignment between their theoretical and observational values.
Cosmological Parameter Symbol Chaotic Planck
@@ 5.2 Cosmological parameters
Table 8: Cosmological parameters obtained from the chaotic inflationary model with a step for $p=1.0004$ compared with the cosmological parameters reported by Planck $2018$ results.
@@
The percentage errors for the age of the universe, matter density, baryon density, dark energy density, and scalar spectral index are $0.007$, $0.16$, $0.04$, $0.09$, and $0.2$, respectively. The effectiveness, accuracy, and precision demonstrated by the chaotic model are again similar to those of the previous ones. It can also be stated that it is slightly better, mainly due to the better precision it achieves when predicting the scalar spectral index.
6 Analysis of the Model Comparison
To assess the validity of different inflationary models with respect to the observed Cosmic Microwave Background (CMB) data, we calculated the $\chi^{2}$ statistic for each model, this was inspired in the analysis made for Ade et al. [37]. The $\chi^{2}$ statistic is given by:
$\chi^{2}=\sum_{i=1}^{n}\frac{(M_{i}-O_{i})^{2}}{\sigma_{i}^{2}},$ (38)
where $M_{i}$ represents the model prediction, $O_{i}$ is the observed data, and $\sigma_{i}$ is the associated uncertainty. To provide a fair comparison across models, we normalized the $\chi^{2}$ by the degrees of freedom $\nu$, which is defined as:
$\nu=N_{\text{data}}-N_{\text{parameters}}.$ (39)
This normalization yields the reduced $\chi^{2}$:
$\chi^{2}_{\nu}=\frac{\chi^{2}}{\nu}.$ (40)
The results for the three inflationary models under consideration are as follows in Table 9.
Inflationary Model $\chi_{\nu}^{2}$ Starobinsky inflationary model 7.52 Generalized Starobinsky inflationary model 0.22 Chaotic inflationary model with a step 0.15
Table 9: $\chi_{\nu}^{2}$ calculated by Starobinsky inflationay model, the generalized Starobinsky inflationary model for $p=1.0004$, the chaotic inflationary model with a step using $c=012$ and $d=0.04$.
The $\chi^{2}_{\nu}$ value for the Starobinsky inflationary model, 7.52, indicates a poor fit to the CMB data, as a $\chi^{2}_{\nu}$ significantly greater than 1 suggests that the model overpredicts the observed values or does not capture the underlying physics accurately. This result implies that the standard Starobinsky model may require further modifications or that it might not be suitable under the current observational constraints.
In contrast, the generalized Starobinsky inflationary model yields a $\chi^{2}_{\nu}=0.22$, and the chaotic inflationary model with a step results in $\chi^{2}_{\nu}=0.15$. Both values are significantly lower than 1, suggesting an overfitting scenario where the models are too flexible relative to the data. These low $\chi^{2}_{\nu}$ values could imply that while these models fit the data well, they may do so at the cost of excessive complexity, capturing noise or fluctuations in the observational data rather than genuine physical features.
Between the two models with low $\chi^{2}_{\nu}$, the chaotic inflationary model with a step exhibits the lowest value, which might be interpreted as slightly better in terms of fitting the observed data. However, such a low $\chi^{2}_{\nu}$ value must be approached with caution, as it may also reflect a lack of robustness rather than true physical relevance.
7 Bayesian Analysis
In order to made a selection of our three best inflationary models also we have perform a Bayesian Analysis using the Cobaya code [25] varying the effective number of neutrino species $N_{\text{eff}}$ an extra $\Delta N_{\text{eff}}$.
After calculate the Bayesian evidence $-\text{log}(Z)$ [38] using the algorithm PolyChord [39] we found the following results:
Inflationary Model $-\text{log}(Z)$ Starobinsky inflationary model $0.803\times 10^{5}\pm 0.384\times 10^{1}$ Generalized Starobinsky inflationary model $0.725\times 10^{5}\pm 0.376\times 10^{1}$ Chaotic inflationary model with a step $0.764\times 10^{5}\pm 0.377\times 10^{1}$
Table 10: The evidence $Z$ is a log–normally distributed, with location and scale parameters $\mu$ and $\sigma$. We denote this as $-\text{log}(Z)=\mu\pm\sigma$.
From table 10, it is necessary to calculate the Bayes ratios $B_{ij}=-\text{log}\left(\frac{Z_{i}}{Z_{j}}\right)$, which are essential to model selection processes [40]. In order to facilitate the calculations let us denote $Z_{1}$, $Z_{2}$, $Z_{3}$ as the Bayesian evidences for the Starobinsky inflationary model, generalized Starobinsky inflationary model, and chaotic inflationary model with a step, respectively.
7.1 Starobinsky inflationary model vs. generalized Starobinsky inflationary model
$B_{21}=\dfrac{0.725}{0.803}\approx 0.903.$ (41)
This means the generalized Starobinsky model is about 0.903 times as likely as the Starobinsky inflationary model.
7.2 Chaotic inflationary model with a step model vs. generalized Starobinsky inflationary model
$B_{23}=\dfrac{0.725}{0.764}\approx 0.949.$ (42)
This means the generalized Starobinsky inflationary model is about 0.949 times as likely as the chaotic inflationary model with a step.
7.3 Starobinsky inflationary model vs. chaotic inflationary model with a step
$B_{31}=\dfrac{0.764}{0.803}\approx 0.951.$ (43)
This means the chaotic inflationary model with a step is about $0.951$ times as likely as the Starobinsky inflationary model.
Analyzing the results above it is possible to conclude that generalized Starobinsky and the chaotic with a step inflationary models are better that the Starobinsky inflationary model. It happens because the Bayes factor corresponds to the relative probability of one model over another [40].
Figure 15: Posterior distributions on the effective number of neutrino species $\Delta N_{\text{eff}}=N_{\text{eff}}-3.046$ for the single–field models discussed in this paper. Starobinsky model inflationary model (red–dotted line), chaotic inflationary model with a step model (green–dotted line), and generalized Starobinsky inflationary model (blue–dotted line).
The previous fact is corroborated by the data presented in Figure 15, where we show the negative logarithm of the posterior distributions of $\Delta N_{\text{eff}}$. It is clear to see that the posteriors of the generalized Starobinsky and chaotic with a step models are similar. Moreover, there is a good amount of data that is near to 00 which means that $\Delta N_{\text{eff}}\approx 0$ or $N_{\text{eff}}\approx 3.046$. The posterior distribution for the Starobinsky model differs from the other distributions which means that other models are more accurate than the Starobinsky model.
8 Constrain of our inflationary models from reheating
In this section we perform an analysis of all our inflationary models considering the number of e-foldings from the end of inflation until the end of reheating, which is given by [41],
$N_{\text{reh}}=\dfrac{4}{1-3\bar{\omega}_{\text{reh}}}\left[61.6-\ln\left(\dfrac{V_{\text{end}}^{\nicefrac{{1}}{{4}}}}{H_{k}}\right)-N_{k}\right],$ (44)
[…]
Following the procedure of Martin et al. [44, 45], and Cook [41], we analyze the dependence of our observables $n_{\mathrm{S}}$ and $r$ by varying the number of e-folds during reheating $N_{\text{reh}}$ with $\bar{\omega}_{\text{reh}}$ for each inflationary model. Then we evaluate these quantities at the duration of reheating $\Delta N\equiv N_{\text{reh}}-N_{e}$. The duration of the reheating period affects the relationship between the scalar spectral index $n_{\mathrm{S}}$ and the scalar–to–tensor ration $r$ and producing different contours.
8.1 Starobinsky inflationary model
For the Starobinsky inflationary model the scalar spectral index $n_{\mathrm{S}}$ and the tensor–to–scalar ratio $r$ in terms of the number of e–folds are calculated from Eqs. (9) and (14) using the slow–roll parameters, and are given by [16],
$n_{\mathrm{S}}$ $\simeq$ $1-\dfrac{2}{\Delta N},$ (45)
$r$ $\simeq$ $\dfrac{12}{\Delta N^{2}}.$ (46)
Fig. 16 shows the contour plot $(n_{\mathrm{S}},r)$ for the Starobinsky inflationary model for three values of $\bar{\omega}_{\text{reh}}=-0.3,-0.2,0$ where we can observed the influence of reheating in our observables.
Figure 16: Contour plot $(n_{\mathrm{S}},r)$ for the Starobinsky inflationay model.
8.2 Generalized Starobinsky inflationary model
For the generalized Starobinsky inflationary model the scalar spectral index $n_{\mathrm{S}}$ and the tensor–to–scalar ratio $r$ in terms of the number of e–folds from the slow–roll parameters, are given by [16],
$n_{\mathrm{S}}$ $\simeq$ $1-\dfrac{8(p-1)\left[C^{2}(p-1)-p(C-1)\right]}{3\left[C(1-2p)+p\right]^{2}},$ (47)
$r$ $\simeq$ $\dfrac{64C^{2}(p-1)^{2}}{3\left[C(1-2p)+p\right]^{2}},$ (48)
where $C\equiv e^{-\frac{4}{3}\Delta N(p-1)}$.
Fig. 17 shows the contour plot $(n_{\mathrm{S}},r)$ for the generalized Starobinsky inflationary model considering three different values of $\bar{\omega}_{\text{reh}}=-0.3,-0.2,0$, and twelve different values for the parameter $p$. In this plot we can observed the influence of the reheating in our observables, and how the generalized Starobinsky inflationary model with $0.995\leq p\leq 1.0004$ reaches the $95\%$ confidence region based on Planck 2018 data when the averaged equation of state during reheating is set to $\bar{\omega}_{\text{reh}}=-0.3$.
Figure 17: Contour plot $(n_{\mathrm{S}},r)$ for the generalized Starobinsky inflationary model for various values of the parameter $p$.
8.3 Chaotic inflationary model with a step
For the chaotic inflationary model with a step the scalar spectral index $n_{\mathrm{S}}$ and the tensor–to–scalar ratio $r$ in terms of the number of e–folds are given by [46],
$n_{\mathrm{S}}$ $\simeq$ $1-\dfrac{2}{\Delta N},$ (49)
$r$ $\simeq$ $\dfrac{8}{\Delta N}.$ (50)
Fig. 18 shows the contour plot $(n_{\mathrm{S}},r)$ for the generalized chaotic inflationary model with a step considering three different values of $\bar{\omega}_{\text{reh}}=-0.3,-0.2,0$, and four set of values of the parameters $c$ and $d$. In this plot we can observed that regardless of the three values of $\bar{\omega}_{\text{reh}}$ , no model enters the $63\%$ confidence region of the Planck 2018 data.
Figure 18: Contour plot of $(n_{\mathrm{S}},r)$ for the chaotic inflationary model with a step for various values of the parameters $c$ and $d$.